# Predictable Drifts in Collective Cultural Attention: Evidence from Nation-Level Library Takeout Data

### v2

@@ 1.3 Contributions from specific books
Although local drift is quite consistent across months, it is unclear to what extent books transition between the different contribution groups (top 100, top 101-1000, etc.) over time. To explore this, we measure the likelihood that a book belonging to a particular contribution group in a given month would remain in the same group or transition to one of the other groups in the following month. We average these transition probabilities over all pairs of consecutive months, and find that only 15% of the 100 books that contribute most to local drift in one month are also part of the 100 highest contributors in the next month (Fig. 3C). Taken together, individual books show substantial variation in their contributions to collective attention drift between consecutive months
@@
To understand which books drive long-term shifts in collective attention, we examine the 1,000 books contributing most to global drift in October 2023 – the last month unaffected by the observed seasonal patterns – and order them by when they reach peak popularity (Fig. 3D). Two patterns emerge. First, rather than distributing evenly across our observation period, these books cluster at the boundaries of our observation window: many peak shortly after the reference month (May 2021), while others peak approaching October 2023. This clustering reveals that global drift arises from two distinct sources: books whose popularity has faded since early in our study period and those gaining recent prominence. Second, the popularity of these high-contributing books is remarkably short-lived. These two patterns – clustering at the edges of the observation period and volatile popularity trajectories – distinguish books driving the global drift from the most popular books and books with highest peaks in popularity (SI Figs. LABEL:popular_overall_counts-LABEL:peak_monthly_popular_counts), revealing that standard measures of book popularity do not fully capture which titles drive long-term collective attention shifts.
@@
In summary, we find individual books to have limited power in explaining drifts in collective attention. Put differently, it is not possible to isolate a few books that account for drifts in a given month, even for drifts in the long term. As such, even if we had perfect predictions of the popularity trajectories of the books with the largest changes in popularity, they would only account for a small percentage of the total drift in collective attention.
1.4 Predicting future drifts from past behavior
@@
So far, we have shown that the collective attention of library loaners is constantly changing. This has a direct impact on the predictability of the system. Further, it implies that the performance of any machine learning model fitted on the cultural consumption at a given point in time, e.g., to recommend books, will deteriorate over time if not updated and retrained. However, since drifts in collective attention exhibit stable patterns, it may be possible to predict attention drift and get an estimate on how fast model performance will decay and how often models should be retrained.
Here, we test the predictability of drifts in collective attention by predicting drifts in 2023 using information about drifts from 2022. As previously observed, drift in collective attention for library loaners is characterized by two components: a steady turnover and seasonal effects. If both of these components are stable across years, it is possible to predict the magnitude of drift in 2023 from the drift observed in 2022. To test this, we make predictions for both local and global drift. When predicting global drift, we use January 2022 and 2023 as baselines, respectively, for quantifying drift in 2022 and 2023. To make predictions, we use the following minimal model: for a given month in 2023, we use the observed (local and global) magnitude of drift for the same period in 2022 as the prediction. For example, if attention has drifted 0.2 bits between two months in 2022, we predict attention to drift with an equal amount between the two months in 2023.
Fig. 4 shows the observed and predicted drifts. Overall, this simple model performs exceptionally well, with minor deviations between the predicted and observed values. Performance remains strong when making predictions for all books (Fig. 4A), or when books have been segmented into different categories (Fig. 4B-D). Further, in SI Fig. LABEL:predictions_subgroups, we report predictions for all combinations of book and loaner subsets and find that the simple model also performs well for all these subsets. This showcases that although it is difficult to predict which items attention will drift towards, predicting the magnitude of drift is possible. This has important implications for how often algorithmic systems that recommend cultural items to users should be updated.
Figure 4: Predictions of future drifts from past drifts. A-D, shows the local and global drifts across all books (A) and for the three main subsets of books B-D, for 2022 and 2023. For each year, global drifts are computed with January as the baseline month.
2 Discussion
@@
For decades, studies of cultural markets have emphasized the difficulty of predicting the ever-changing attention of cultural consumers [41, 42, 17, 18, 43]. We show that changes in the collective attention of Danish book loaners are consistently high and cannot be attributed to only a few books. Further, our study shows that the collective redistribution of attention across books in Danish libraries has identifiable patterns that can be predicted with a high degree of accuracy, even for specific groups of products and audiences. While we examine only broad product and audience groups, our findings suggest that predictable collective attention dynamics could exist for more fine-grained groupings, such as specific sub-genres.
@@
Our findings raise important questions about the mechanisms underlying collective cultural change, specifically processes of individual choice and social influence. Relationships between individual choice, social influence, and cultural change have previously mainly been investigated in simulation studies [44, 45, 46, 47] and experimental work [17, 18, 48, 49, 50, 51]. Detailed registry data on individual loaners’ choices, such as the one we used in this study, can open new opportunities to study relationships between behavior at the individual-, group-, and macro-levels. A well-developed understanding of these mechanisms could be used to explain differences in how collective cultural attention evolves for different types of products and groups of consumers.
Our dataset contains comprehensive records of all Danish public library loans; however, our findings are not necessarily guaranteed to generalize across all book consumption. Book consumption is most likely different when there is an associated economic cost. For example, since borrowing books at the library is free, users may explore titles more widely than when they have to buy new books in a bookstore. Similarly, libraries feature a wide selection of both new and older titles, whereas bookstores prioritize newer releases and popular titles. These factors likely influence the drift of collective attention, but investigating their impact is beyond the scope of this work. Additionally, although we aggregate various media formats (physical books, ebooks, and audiobooks) to maximize coverage, consumption patterns may differ across these formats. Future work should look into this.
@@
The extent to which our results can be generalized to other cultural markets is an empirical question; however, we expect similar forms of collective attention dynamics to apply widely. For instance, other cultural markets display variations of the drift components we observe for library loans: a stable turnover in consumer preferences and season-specific item consumption [52, 53]. Nonetheless, the importance of individual products for drift dynamics could differ significantly between markets. In markets with fewer products and more unequal distributions of popularity, a small number of products could account for a larger proportion of drift. In such markets, shocks by individual products could disrupt otherwise stable drift patterns, causing more unpredictable drift dynamics than what we observe for library loans. We hope that future studies will explore how characteristics of drifts in collective attention compare across cultural markets.
@@
Recommender systems are another potential source of variation in consumer attention dynamics between markets. Recommender systems are pervasive in many current cultural markets, and their effects on drifts in collective consumer attention should be studied more closely. However, because these systems are so deeply embedded in most real cultural markets, isolating their effects is a challenge [54]. Even public libraries rely heavily on centralized recommendation systems, both on their websites for book search and reservations, and via librarians’ curation of books in each local library.
@@
Our findings have major implications for adapting to changing consumer attention. With reliable expectations of drifts in consumer attention, individuals and organizations interested in predicting and modeling cultural consumption can make more informed decisions about when machine learning models require updating. As we show, drifts in collective attention can be faster for certain product types and consumer groups (Fig. 2). Model updating should take into account these group-specific drift differences, as failing to do so could result in problematic biases towards groups with faster drifts in collective attention. For example, drift is higher for females and for 65+ year demographics, meaning models need to be updated more frequently to accurately reflect their tastes. The consequences of drifts in collective attention will depend on the specific model. In particular, many recommender systems are based on collaborative filtering approaches that capture associations between users and items [55]. In these models, performance declines can result not only from drift in item popularity but also from drift in the user base and drift in the association between users and items. Our study focuses solely on item popularity, leaving open important questions about the predictability of drift in other data used to model cultural consumption.
@@
We envision the methods developed in this study to be used to study changes in collective behavior in other domains outside of cultural markets. Like cultural consumption, behaviors that are socially driven are generally difficult to predict, including information cascades, social influence, and social interactions [56, 57, 58, 59]. For such many behaviors, it may be possible to make higher-order predictions of behavior that can be used to contextualize and adapt to uncertain predictions of individual phenomena like the popularity of specific topics or items.
3 Methods
@@
Data Access & Anonymity. All data processing and analysis was conducted in-house at Statistics Denmark, which is the national statistical agency of the country, in accordance with their data protection protocols. This means that all individual-level data, both the library takeout dataset and the sociodemographic loaner information, are pseudonymized and kept strictly confidential. Due to the sensitive nature of the data, it is not possible to openly share the data, but it is possible to apply for access to the data through Statistics Denmark.
@@
The Danish Nation-Scale Library Takeout Dataset. The library dataset used in this study is based on administrative registers of library takeouts from Danish public libraries, collected by Statistics Denmark. The dataset is continuously updated, and the version used in this study spans from January 1, 2020, to December 31, 2023, and includes all loans from public libraries, including Ereolen (the Danish national platform for digital books), but excluding university and school libraries.
@@
Sociodemographic Loaner Information. All variables except sex are dynamic in our dataset. This means that loaners can belong to different subgroups at different points in time. Below, we describe how we derived the different loaner categories. Age is computed directly from each loaner’s birthdate at the time of each loan. Residence Area is based on standard definitions provided by Statistics Denmark, which divides municipalities into five types (https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/kommunegrupper), which we collapse into two main types: Large city municipality, which is defined by a municipality with above 100,000 citizens or has a high job availability (see https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/kommunegrupper for specifics), and Town or rural municiplaity, which covers all other municipalities. Level of Education is recorded as the loaner’s maximum attained education level at the time of each loan. The three categories are based on the DISCED-15 classification (see https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/disced15-udd), which groups education into 15 main categories.
Library Data Preprocessing.
@@
Item Filtering. Around 10 million loans in the dataset refer to non-books (see SI Fig. LABEL:material_type_distribution for the distribution of loans across material types), which were removed from analysis.
@@
Book Identification. Each item is associated with a unique identification number, known as a faust number. Because each edition of a book gets a separate faust number, one title can have several associated faust numbers. Since we are interested in the evolution in the popularity of titles, regardless of the medium (physical book, audiobook, etc.), we treat different media variants of the same book as a single entity. To derive a unique identifier independent of the edition and medium, the following heuristic algorithm was applied:
@@
This process reduced the unique number of items from approximately 784,000 to around 660,000.
Book Category Book categories (adult fiction, adult nonfiction, children’s) were provided by the Danish Bibliographic Centre and not further preprocessed.
Measuring Changes in Collective Consumer Attention.
Jensen-Shannon Divergence and Book-Level Contributions. In each month, we compute the popularity (number of loans) of each book $i$ relative to the total amount of loans in that month. The drift in collective attention between those months is the JSD between the distributions of popularity in two time months $P$ and $Q$, which can be expressed as a linear combination of changes to the popularity of each individual book:
$\text{JSD}(P,Q)$ $=\sum_{i}\Bigl(\overbrace{{p_{i}\log\frac{2p_{i}}{p_{i}+q_{i}}}+{q_{i}\log\frac{2q_{i}}{p_{i}+q_{i}}}}^{\text{Partial JSD}_{i}}\Bigr)$ (1)
where $p_{i}$ is the relative popularity of book $i$ in the first month, $q_{i}$ is the relative popularity of book $i$ in the second period. The partial JSD represents the contribution of each book to the overall JSD. If there is no change to a book’s relative popularity, it does not contribute to the JSD. For any measure of drift, the relative contribution of a group of books to that drift is simply calculated as the sum of partial JSDs for books in that group divided by the JSD.
Estimation of Jensen-Shanon Divergence. Empirical estimation of information-theoretic quantities from samples is known to suffer from bias, particularly when the number of categories (here: unique books) is larger than the total number of sampled counts (here: loans) [60]. To maintain a consistent relation between changes in overall preferences and changes in preferences for individual books, we use the maximum likelihood estimator when computing JSDs for all loaners.
@@
4 Acknowledgements
We are grateful for the help with data collection by the Data Science Lab at Statistics Denmark and Ea Hoppe Blaabæk.

### v3

@@ Results
Although each group’s contribution to local drift is relatively stable across months, it is unclear to what extent books transition between the different contribution groups (top 100, top 101-1,000, etc.) over time. To explore this, we measure the likelihood that a book belonging to a particular contribution group in a given month would remain in the same group or transition to one of the other groups in the following month. We average these transition probabilities over all pairs of consecutive months, and find that only 14% of the 100 books that contribute most to local drift in one month are also part of the 100 highest contributors in the next month (Fig. 3C). This means that a very large shift in a book’s popularity are typically followed by a smaller shift. Taken together, individual books show substantial variation in their contributions to drift of collective attention between consecutive months.
@@
To understand which books drive long-term shifts in collective attention, we examine the 1,000 books contributing most to global drift in October 2024 – the last month in the data unaffected by the observed seasonal patterns – and order them by when they reach peak popularity (Fig. 3D). Two patterns emerge. First, rather than distributing evenly across our observation period, the peak popularity of these books cluster at the boundaries of our observation window: many peak shortly after the reference month (May 2021), while others peak approaching October 2024. This ordering reveals that global drift arises from two distinct sources: books whose popularity has faded since early in our study period and those gaining recent prominence. Second, the popularity of these high-contributing books is remarkably ephemeral. These two patterns – clustering at the edges of the observation period and volatile popularity trajectories – distinguish books driving the global drift from the most popular books and books with highest peaks in popularity (Supplementary Figs. S26-27), revealing that standard measures of book popularity do not fully capture which titles drive long-term collective attention shifts. In Supplementary Fig. S28, we show that the seasonal drift, identified above, are largely driven by specific Christmas-themed children’s titles which regain popularity each holiday season but remain relatively ‘unpopular’ during the rest of the year.
@@
In summary, we find individual books to have limited power for explaining drift in collective attention. As such, even if we could perfectly predict the popularity trajectories of the books with the largest changes in popularity, they would only account for a small percentage of the total drift in collective attention. This underscores the importance of considering the full ‘cultural system’ and not just individual product trajectories when analyzing patterns of consumer attention.
Generalizability and predictability of drift patterns
@@
So far, we have shown that the collective attention of library loaners is constantly changing. This has a direct impact on the predictability of the system. Further, it implies that the performance of any machine learning model fitted on the cultural consumption at a given point in time, e.g., to recommend books, will deteriorate over time if not continually kept-up-to-date and retrained. However, since drift in collective attention exhibits stable patterns, it may be possible to predict attention drift and get an estimate on how fast model performance will decay and how often models should be retrained.
To examine whether drift patterns are stable over time, we quantify drift between all pairs of months over the five-year period from 2020 to 2024, excluding months affected by COVID-19 lockdowns. A visual inspection of drifts in Fig. 4A shows that the patterns we previously identified, for a subset of months, are remarkably consistent across all months; drift in collective attention generally increases with distance in time between any two months, and identical seasonal patterns appear across all years. In the SI, we show that the drift patterns are equally consistent for book subcategories and that the pattern of increasing drift over time also holds during COVID-19 lockdown months, although drift in these months are generally higher (Supplementary Figs. S29-33). In Fig. 4B–F, we further show that both local and global drift remain highly stable from year to year, for all books and across the different book categories included in our analysis. This stability suggests that future changes in collective attention can be approximated from past changes.
Figure 4: Predictability of collective attention drift across years. A shows all pairwise drifts between months from 2020 to 2024 (excluding COVID-affected months), illustrating the increase in drift with temporal distance and the consistency of seasonal patterns. B shows the local and global drift for all books, and C–F show the corresponding drift for the major book subcategories (GA stands for General Audience). For each year, global drift is computed using January as the baseline month. Since January 2021 was affected by a partial COVID-19 lockdown, global drift is not computed in 2021.
We now consider whether drifts in future years can be predicted using observations from earlier years. For this purpose, we use a minimal model. For any pair of months in a given year, we use the drift observed for the same period in the previous year as the prediction. For example, if attention drifted by 0.2 bits between two months in 2022, we predict that the attention between the same two months in 2023 will also drift by 0.2 bits. Despite its simplicity, this model performs remarkably well. Across all loaner subgroups, the mean absolute percentage error (MAPE) for predicting one year’s drift from the preceding year remains below 10% with slightly lower errors for general audience books than children’s books (Supplementary Figs. S34-38). Prediction accuracy degrades when using older reference years for making predictions, suggesting subtle long-term changes in drift patterns.
Together, these results show that, although we cannot predict which specific items will gain or lose attention, we can reliably predict the magnitude of drift in collective attention. This has important implications for policy makers, regulators, and practitioners dealing with how often algorithmic systems that recommend cultural items to users should be updated and evaluated.
Discussion
@@
For decades, studies of cultural markets have emphasized the difficulty of predicting the ever-changing attention of cultural consumers [41, 42, 12, 13, 43]. We show that changes in the collective-level attention of Danish book loaners are consistently high and cannot be attributed to only a subset of titles. However, our study shows that the collective redistribution of attention across books in Danish libraries has identifiable global patterns that can be predicted with a high degree of accuracy, even for specific groups of products and audiences. While we examine only broad product and audience groups, our findings suggest that predictable collective attention dynamics could exist for more fine-grained groupings, such as sub-genres.
@@
Our findings raise important questions about the mechanisms underlying collective cultural change, specifically processes of individual choice and social influence. Relationships between individual choice, social influence, and cultural change have previously mainly been investigated in simulation studies [44, 45, 46, 47] and experimental work [12, 13, 48, 49, 50, 51]. Detailed registry data on individual loaners’ choices, such as the one we used in this study, can open new opportunities to study relationships between behavior at the individual-, group-, and macro-levels, especially when combined with administrative data on the social opportunity structures of each individual loaner [52].
Inferring mechanisms from purely observational data, however, requires caution. Although simple mechanistic models – such as those based on collective memory [22], neutral drift [53], and information cascades [54] – have been shown to reproduce patterns of cultural change in empirical data, such patterns are often compatible with multiple alternative explanations [55, 56, 57]. We have avoided imposing any particular model of individual-level selection and instead focused on the patterns in the collective attention that emerge from loaners’ aggregated choices. We have shown that these patterns are both interpretable and predictable without knowing the underlying mechanisms. At the same time, they place constraints on which mechanisms could plausibly generate them. For example, the recurrence of holiday-themed children’s books indicates that models which do not account for fluctuating selections cannot reproduce the observed dynamics. That such fluctuations are observed for children’s books but not general audience books further suggests heterogeneity in the mechanisms driving selection across different book types. Heterogeneity of selection mechanisms is incompatible with models based on pure reinforcement like neutral models, which have been used extensively for to explain cultural change [58, 53].
@@
The extent to which our results can be generalized to other cultural markets is an empirical question, however, we expect similar forms of collective attention dynamics to apply widely. For instance, other cultural markets display variations of the drift components we observe for library loans: a stable turnover in consumer preferences and season-specific item consumption [53, 59]. Nonetheless, the importance of individual products for drift dynamics could differ significantly between markets. In markets with fewer products and more unequal distributions of popularity, a small number of products may account for a larger proportion of drift. In such markets, shocks by individual products could disrupt otherwise stable drift patterns, causing more unpredictable drift dynamics than what we observe for library loans. We hope that future studies will explore how characteristics of drifts in collective attention compare across various cultural markets.
Our dataset contains comprehensive records of all Danish public library loans; however, our findings are not guaranteed to generalize across all book consumption. Book consumption is most likely different when there is an associated economic cost. For example, borrowing books at the library is free which may cause users to explore titles more widely compared to when they have to buy books in a bookstore. Similarly, libraries feature a wide selection of both new and older titles, whereas bookstores prioritize newer releases and popular titles. These factors could likely influence the drift of collective attention.
@@
Recommender systems are another potential source of variation in consumer attention dynamics between markets. Recommender systems are pervasive in many current cultural markets, and their effects on drift in collective consumer attention should be studied more closely. However, because these systems are so deeply embedded in most real cultural markets, isolating their effects is a challenge [60]. Even public libraries rely heavily on centralized recommendation systems, both on their websites for book search and reservations, and via librarians’ curation of books in each local library.
@@
Our findings have major implications for adapting to changing consumer attention. With reliable expectations of drifts in consumer attention, individuals and organizations interested in predicting and modeling cultural consumption can make more informed assessments of when machine learning models require updating. As we show, drifts in collective attention can be faster for certain product types and consumer groups (Fig. 2). Model updating should take these group-specific drift differences into account. Failing to do so could result in problematic biases towards groups with faster drift in collective attention. For example, in our data, drift is higher for women and for 65+ year-old demographics, meaning models need to be updated more frequently to accurately reflect their tastes. The consequences of drift in collective attention will depend on the specific model. In particular, many recommender systems capture associations between users and items [61]. In these models, reductions in performance can result not only from drift in item popularity but also from drifts in the user base and drifts in the associations between users and items. Our study focuses solely on item popularity, leaving open important questions about the predictability of drift in other data used to model cultural consumption.
@@
We envision the methods developed in this study to be used to study changes in collective behavior in other domains outside of cultural markets. Like cultural consumption, behaviors that are socially driven are generally difficult to predict, including information cascades, social influence, and social interactions [62, 63, 54, 64]. For such behaviors, it may be possible to identify and predict stable macro-level patterns of changing behavior that can be used to contextualize and adapt to uncertain predictions of individual phenomena, like the popularity of specific topics or items.
Methods
Ethics and anonymity
@@ Methods
Since this study was based on a secondary analysis of administrative registers, no ethical approval was required for this study. All data processing and analysis was conducted in-house at Statistics Denmark, the national statistical authority of Denmark, in compliance with their data protection protocols and the requirements of the Danish Data Protection Agency. This means that all individual-level data, both the library takeout dataset and the sociodemographic loaner information, were pseudonymized and kept strictly confidential.
Data Access
Due to the sensitive nature of the data, it is not possible to openly share the data, but it is possible to apply for access to the data through Statistics Denmark.
The Danish Nation-Scale Library Takeout Dataset
@@
The library dataset used in this study is based on administrative registers of library takeouts from Danish public libraries, collected by Statistics Denmark. The dataset is continuously updated, and the version used in this study spans from January 1, 2020, to December 31, 2024, and includes all loans from public libraries, including Ereolen (the Danish national platform for digital books), but excluding university and school libraries.
Sociodemographic Loaner Information
@@
All variables except sex are dynamic in our dataset. This means that loaners can belong to different subgroups at different points in time. Below, we describe how we derived the different loaner categories. Age is computed directly from each loaner’s birthdate at the time of each loan. Residence Area is based on standard definitions provided by Statistics Denmark, which divides municipalities into five types (https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/kommunegrupper), which we collapse into two main types: 1) Large city municipality, which is defined by a municipality with above 100,000 citizens or has a high job availability (see https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/kommunegrupper for specifics), and 2) Town or rural municiplaity, which covers all other municipalities. Level of Education is recorded as the loaner’s maximum attained education level at the time of each loan. The three categories are based on the DISCED-15 classification (see https://www.dst.dk/en/Statistik/dokumentation/nomenklaturer/disced15-udd), which groups education into 15 main categories.
Library Data Preprocessing
Item Filtering
@@
Around 12 million loans in the dataset refer to non-books (see Supplementary Fig. S2 for the distribution of loans across material types), which were removed from analysis.
Book Identification
@@
Each item is associated with a unique identification number, known as a faust number. Because each edition of a book gets a separate faust number, one title can have several associated faust numbers. Since we are interested in the evolution in the popularity of titles, regardless of the medium (physical book, audiobook, etc.), we treat different media variants of the same title as a single entity. To derive a unique identifier independent of the edition and medium, the following heuristic algorithm was applied:
@@
This process reduced the unique number of items from approximately 1,060,000 to around 760,000.
Book Category
Book categories (general audience fiction, general audience nonfiction, children’s) were provided by the Danish Bibliographic Centre.
Measuring Changes in Collective Consumer Attention
Jensen-Shannon Divergence and Book-Level Contributions
In each month, we compute the popularity (number of loans) of each book $i$ relative to the total amount of loans in that month. The drift in collective attention between those months is the JSD between the distributions of popularity in two months $P$ and $Q$, which can be expressed as a sum of individual book contributions:
$\text{JSD}(P,Q)$ $=\sum_{i}\Bigl(\overbrace{{0.5p_{i}\log\frac{2p_{i}}{p_{i}+q_{i}}}+{0.5q_{i}\log\frac{2q_{i}}{p_{i}+q_{i}}}}^{\text{Partial JSD}_{i}}\Bigr)$
where $p_{i}$ is the relative popularity of book $i$ in the first month, $q_{i}$ is the relative popularity of book $i$ in the second month. The partial JSD represents the contribution of each book to the overall JSD. If there is no change to a book’s relative popularity, it does not contribute to the JSD. For any measure of drift, the relative contribution of single book to that drift is simply calculated as the partial JSD of that book divided by the total JSD.
Estimation of Jensen-Shanon Divergence
Empirical estimation of information-theoretic quantities from samples is known to suffer from bias, particularly when the number of categories (here: unique books) is larger than the total number of sampled counts (here: loans) [65]. To maintain a consistent relation between changes in overall collective attention and changes in collective attention for individual books, we use the maximum likelihood estimator when computing JSDs for all loaners.
@@
[…]
Acknowledgements We are grateful for the help with data collection by the Data Science Lab at Statistics Denmark and Ea Hoppe Blaabæk.
Competing interests
The authors declare no competing interests.
Author contribution
Both authors designed the study, WL performed the analysis, both authors analyzed the data, and wrote the manuscript.